### v1

@@ Abstract
Control-flow attestation (CFA) unifies the worlds of control-flow integrity and platform attestation by measuring and reporting a target’s run-time behaviour to a verifier. Trust assurances in the target are provided by testing whether its execution follows an authorised control-flow path. The problem has been explored in various settings, such as assessing the trustworthiness of cyber-physical systems, Internet of Things devices, cloud platforms, and many others. Despite a significant number of proposals being made in recent years, the area remains fragmented, addressing different adversarial behaviours, verification paradigms, and deployment challenges. In this paper, we present the first survey of control-flow attestation, examining the core ideas and solutions in state-of-the-art schemes. In total, we survey over 30 papers published between 2016–2024, consolidate and compare their key features, and pose several challenges and recommendations for future research in the area.
@@ 1 Introduction
Today’s computing platforms are susceptible to attacks manipulate software behaviour at run time. Control-flow attacks, such as return-oriented programming (ROP), have become commonplace that execute unintended program code paths in order to perform unauthorised functions using existing code fragments. These fragments are combined in sequence in a way that evades conventional defences, such as code signing and other static integrity-checking methods. The challenge of run-time threats has underpinned the area of control-flow integrity (CFI), which aims to ensure that programs adhere to authorised paths of execution [1]. This can be achieved, for example, by verifying the legitimacy of the destination address of every control-flow transfer (e.g. jump and call instructions). Other mechanisms have also been developed, such as code pointer integrity (CPI) [2], which segregates pointers from other data in memory; and data execution prevention (DEP), which prevents code from being executed from regions of memory that are not explicitly marked as executable.
@@
  • Using a common set of criteria, we compare existing CFA schemes with respect to their trust assumptions, target control-flow transfers, attack coverage, and other salient features in order to elucidate their features, advantages, and shortcomings.
@@ 1.1 Related Work
In some of these surveys [16, 18, 21], control-flow attestation is described briefly but not comprehensively, considering only early control-flow attestation schemes (e.g. [4, 8]). They do not present a dedicated treatment to the large number of CFA proposals that have emerged over the past decade. We directly address this gap in this paper. Survey papers on control-flow integrity more generally have also been published on verification methods [22] and hardware-based CFI systems [23]. We emphasise that CFI is a related concept necessary for understanding CFA schemes, but it is not directly concerned with attesting target platforms.
This paper offers a comprehensive treatment of proposals that attest devices using run-time control-flow behaviours. Specifically, we consider proposals that contain mechanisms for securely measuring and reporting the state of a proving device. Proposals that focus on detecting only control-flow integrity violations are considered useful supplementary literature, but they are orthogonal to this work.
@@ 2 Fundamentals of Control-Flow Integrity
When the CFI security property [1] holds for a given program, $P$, we say that $P$ adheres to an authorised walk, $w=\{v_{0},v_{1},\dots v_{n}\}$, for a connected sequence of vertices with starting and final vertices, $v_{0}$ and $v_{n}$ respectively. [^1] Existing literature often refers to the term ‘path’, even though this can be considered a misnomer in graph theory. A path is a trail in which all vertices (and thus all edges) are distinct, whereas a walk is more general: it is a finite or infinite sequence of edges which joins a sequence of vertices. For a given program, the implementation of control-flow integrity for a desired program is achieved generally in two phases:
@@
    CFI enforcement: This is performed at runtime and involves a reference monitor that checks and enforces the program’s adherence to the CFG derived in the previous step. A common CFI enforcement method is through program instrumentation [1]. Here, a checkpoint $C(p,t)$ is inserted at a control transfer instruction at point $p$ in BBL, with a target, $t$. Here, $p$ and $t$ represent a logical CFG edge between one BBL $v_{i}$ and a target BBL, $v_{j}$. Before executing any control transfer instruction, the program passes control to the reference monitor that verifies whether $C(p,t)$ is associated with $(v_{i},v_{j})\in E$. That is, the monitor ensures that the transfer complies with the CFG. If the check fails, the monitor may half the program’s execution, throw an exception, or take another enforcement action.
@@ 2.1.1 Control Data Attacks
The primary class are attacks that directly modify control data used to that directly influences the processor’s program counter (PC) during execution by modifying return addresses and function pointers [25]. A fundamental attack vector, which we refer to as code injection, involves inserting malicious code into executable memory and redirecting the PC to that code.
@@
Stack-based buffer overflows are a canonical example, which involves overwriting function return addresses and other variables on the stack to redirect execution to code to a malicious payload (e.g. shellcode). Heap-based attacks are another approach, involving overflowing buffers on the heap and affecting adjacent heap metadata and objects (e.g. function pointers) that redirect control flow.
@@
A common example of a code-reuse attack is return-oriented programming (ROP) [26], which relies on return instructions to chain together a sequence of gadgets. Each gadget typically concludes with a return instruction; the next gadget in the sequence starts based on the tampered return address of the previous one. It is widely known ROP attacks can achieve Turing-completeness [27]. Jump- (JOP [28]) and call-oriented programming (COP [29, 30]) attacks are two other code reuse attacks, which involve tampering forward control-flow transfers, namely indirect jumps and indirect calls respectively, in order to chain sequences of gadgets.
@@ 2.1.2 Non-Control Data Attacks
Rather than directly modifying control data, e.g. function pointers, non-control data attacks rely on the manipulation of more general data to indirectly change a program’s control flow [31, 32, 33]. Detecting non-control-data attacks is more challenging because they technically traverse legitimate edges of the program’s CFG, even though malicious functionality ensues as a consequence. Let us consider a motivating example. An adversary may execute arbitrary commands on a Linux system if they can influence the arguments to a execve() call in the target program. The target program’s CFG is respected, even though the resulting behaviour is unauthorised. Data attacks can be realised in the following ways:
@@
Branch variable attacks are an example of (A.3) in which an adversary influences data used in determining the outcome of conditional branches, i.e. in ‘if’ and ‘switch’ statements. The resulting data causes another legitimate branch to be taken, which nevertheless results in unintended behaviour. [^3] It is important to recognise this semantic gap: just because legitimate edges of a CFG are followed does not mean that unintended or unauthorised behaviour more abstractly can be performed. In a similar case, the adversary may attempt to manipulate the data that is used in the termininating condition of loops, e.g. ‘while’ and ‘for’ statements. Such attacks may also involve manipulating the data of loop termination conditions, which we refer to as a ‘loop variable attack’.
@@
    Data-oriented attacks: Executes a sequence of existing data gadgets on attacker-controlled inputs in order to perform malicious actions while respecting the target program’s CFG.
@@ 2.2 Control-Flow Transfer Types
In CFI and CFA, monitoring different control-flow transfer types is used to detect particular unauthorised paths within a target program. We note that it is often undesirable to monitor every possible control-flow due to the significant performance overhead that this can occur, especially on constrained devices. Indeed, not all control-flow transfers are relevant to the security of a system, and it is intuitive to monitor transfers that are common targets for particular exploits of interest. For example, one may wish to pay particular attention to indirect jumps and calls in order to address ROP and JOP attacks. In Figure 2, we give some examples of such transfers used by X86-64 platforms. We now describe the control-flow transfer types have been used in existing work:
@@
Figure 2: X86-64 assembly examples of various control-flow transfer types.
@@
    All control-flow transfers: This is the most expansive definition of $\mathcal{C}$. We assume that $\mathcal{C}$ is a maximal set, or has the greatest cardinality, for the target architecture. In practice, all possible control-flow instructions are examined—conditional jumps, direct calls, indirect calls, etc.—imposing the greatest overhead during monitoring and measurement. It is also the most invasive, requiring the heaviest degree of instrumentation of the target program (see §6). Monitoring all control-flow transfers has been used in many existing proposals, with a heavy focus on proposals requiring dedicated hardware detection circuits [37, 12, 34, 38, 36, 39, 40, 9, 41].
@@
    Conditional direct branches/jumps: Includes all branching and jump instructions, represented as assembly or machine code, that alter the execution sequence based on a specific condition. If the condition is true, then the program counter is directly set to a specified address, continuing execution from that address. In direct jumps, the target address or label is explicitly stated in the instruction. The address is fixed at compile time and does not change at runtime.
@@
    Return instructions: Return instructions are used at the end of a function to return to the location of the PC before the function was invoked. On X86, the instruction ret pops the top value from the stack, which should be the return address pushed onto the stack by a previous call instruction, before jumping to that address.
@@ 3 Introducing Control-Flow Attestation
$\mathcal{C}$ denotes the control-flow transfer types (e.g. direct jumps, indirect calls) being monitored, which were presented in §2.2. A general requirement is that CFA scheme must be able to monitor unique information about a BBL associated with a single control-flow transfer type. During collection, this information may include the addresses of a BBL’s first and final instructions; the control flow transfer’s source and destination addresses, which delineates different BBLs; a unique label identifying the BBL; [^5] IDs are used to overcome issues with Address Space Layout Randomisation (ASLR) in which instructions’ memory addresses change with each program execution. Using only instruction addresses, for example, cannot delineate between different BBLs in this scenario. To solve this challenge, the $T_{\mathcal{P}}$ is instrumented to insert BBL IDs to mark the beginning of BBLs, which do not exist in the original target program. and the type of control-flow transfer and its complete instruction representation, e.g. as machine code.
@@
In Figure 3, we illustrate the knowledge areas that characterise their core elements. We posit that schemes can be delineated according to their attestation paradigm; the trust anchor; the instrumentation method for collecting measurements; the measurement method; and, finally, the adversarial model under consideration. The coming sections describe the proposals and applications within each area.
@@ 4 Prover-Verifier Paradigms
We recognise that the majority of CFA schemes operate under an interactive paradigm, which is initiated and verified by a (remote) verifier using a challenge-response protocol. This follows a conventional model of remote attestation used by trusted computing systems, such as the Trusted Platform Module (TPM). However, other approaches have also been developed in recent years.
4.1 Conventional Attestation
@@ 4.2 Continuous Reporting
Rather than reporting measurements in the context of a challenge-response protocol, LiteHAX by Dessouky et al. [34] reports measurements continuously $\mathcal{P}$to $\mathcal{V}$. In this work, there is an initial offline phase where $\mathcal{V}$ generates the CFG of $T_{\mathcal{P}}$. However, during the online phase, $\mathcal{P}$ sends attestation reports to $\mathcal{V}$ in a one-way, continuous manner. There is no formal challenge-response protocol at run-time. We illustrate this in Figure 5. ACFA [44] similarly reports measurements $\mathcal{P}$to $\mathcal{V}$. The transmission of the attestation report is triggered by one of three conditions: (1) the expiration of a timer, which enables $\mathcal{P}$ to send reports periodically; (2) the memory designated for storing measurement results becomes full; or (3) the completion of the attested code execution. Another difference with ACFA is that, after sending the attestation report, $\mathcal{P}$ pauses the target program’s execution until $\mathcal{V}$ returns a response to notify whether execution can be resumed.
Figure 5: LightHAX: a uni-directional CFA scheme (from [34]).
@@ 4.3 Hybrid Attestation
We define hybrid attestation as a combination of local monitoring and remote attestation. Geden and Rasmussen [47] follows the same offline stage as conventional remote attestation where $\mathcal{V}$ extracts a CFG, known as a runtime integrity model, before deploying the proving device. The model is stored internally in a discrete hardware security module (HSM) aboard $\mathcal{P}$. The HSM continuously takes control-flow events as input from the CPU in real-time in order to thwart time-of-check-to-time-of-use (TOCTOU) problem in conventional remote attestation. A remote verifier can acquire a report from the HSM, showing whether any security policy violations were detected over the course of the target program’s execution.
@@
GuaranTEE by Morbitzer et al. [10] investigates control-flow attestation within untrusted cloud environments. The work differs in that $\mathcal{P}$ and a local verifier, $\mathcal{V}^{\prime}$, run on the same cloud-based virtual machine (VM) in Intel SGX enclaves. The $\mathcal{P}$ is assumed to host a security-sensitive application that may be targeted by control-flow attackers; a signing service for Covid-19 digital passports is given as a motivating example. $\mathcal{V}^{\prime}$ contains the logic for measuring and analysing valid control-flow transfers, which are passed through a trampoline implemented in $\mathcal{P}$. We illustrate this in Figure 6. The target program’s instrumentation is performed through extensions to the LLVM compiler framework, allowing $\mathcal{P}$’s full CFG to be constructed pre-deployment. GuaranTEE supports a form of remote and local attestation. At launch-time, each SGX enclave mutually attests the other on the local machine, while enabling a remote $\mathcal{V}$ to retrieve attestation reports from $\mathcal{V}^{\prime}$.
@@ 4.5 Collective Attestation
Figure 7: The ARCADIS system model (from [41]).
@@
Li et al. [49] proposed CFRV that tackles attestation in networks of peer-to-peer IoT devices, where devices may act as $\mathcal{P}$ and $\mathcal{V}$ in a given session. The authors address the problem of adversarial verifiers who submit malicious attestation challenges that purposely induce a particular control-flow path resulting in unintended behaviour. To address this, the authors combine secret sharing and mutual attestation in order to create an agreed attestation measurement between multiple devices.
@@
Figure 8: System overview of DIAT (from [45]). Unlike conventional schemes, measurements from each sub-system’s CFGs are aggregated within attestation reports.

### v3

@@ Abstract
Control-flow attestation unifies the worlds of control-flow integrity and platform attestation by measuring and reporting a target’s run-time behaviour to a verifier. Trust assurances in the target are provided by testing whether its execution follows an authorised control-flow path. The problem has been explored in various settings, such as assessing the trustworthiness of cloud platforms, cyber-physical systems, and Internet of Things devices. Despite a significant number of proposals being made in recent years, the area remains fragmented, addressing different adversarial behaviours, verification paradigms, and deployment challenges. In this paper, we present the first survey of control-flow attestation, examining the core ideas and solutions in state-of-the-art schemes. In total, we survey over 30 papers published between 2016–2024, consolidate and compare their key features, and pose several challenges and recommendations for future research in the area.
@@ 1 Introduction
Today’s computing platforms are susceptible to attacks which manipulate software behaviour at run time. Control-flow attacks, such as return-oriented programming (ROP), have become commonplace that execute unintended program code paths in order to perform unauthorised functions using existing code fragments. These fragments are combined in sequence in a way that evades conventional defences, such as code signing and other static integrity-checking methods. The challenge of run-time threats has underpinned the area of control-flow integrity (CFI), which aims to ensure that programs adhere to authorised paths of execution [1]. This can be achieved, for example, by verifying the legitimacy of the destination address of every control-flow transfer (e.g. jump and call instructions). Other mechanisms have also been developed, such as code pointer integrity (CPI) [2], which segregates pointers from other data in memory; and data execution prevention (DEP), which prevents code from being executed from regions of memory that are not explicitly marked as executable.
@@
  • Using a common set of criteria, we compare existing CFA schemes with respect to their methods, trust assumptions, target control-flow transfers, and adversarial models in order to elucidate their features, advantages, and shortcomings.
@@ 1.1 Related Work
In some of these surveys [16, 18, 21], control-flow attestation is described briefly but not comprehensively, considering only early control-flow attestation schemes (e.g. [4, 8]). They do not present a dedicated treatment to the large number of CFA proposals that have emerged over the past decade. We directly address this gap in this paper. Survey papers on control-flow integrity more generally have also been published on verification methods [22] and hardware-based CFI systems [23]. This paper offers a comprehensive treatment of proposals that attest devices using run-time control-flow behaviours. Specifically, we consider proposals that contain mechanisms for securely measuring and reporting the state of a proving device. Proposals that focus on control-flow integrity, but do not provide attestation, are considered useful supplementary literature; however, they are orthogonal to this work.
@@ 2 Fundamentals of Control-Flow Integrity
When the CFI security property [1] holds for a given program, $P$, we say that $P$ adheres to an authorised walk, $w=\{v_{0},v_{1},\dots v_{n}\}$, for a connected sequence of vertices with starting and final vertices, $v_{0}$ and $v_{n}$ respectively. [^1] Existing work often uses the term ‘path’ instead of ‘walk’, even though this can be considered a misnomer in graph theory. A path is a trail in which all vertices (and thus all edges) are distinct, whereas a walk is more general: it is a finite or infinite sequence of edges which joins a sequence of vertices. For a given program, the implementation of control-flow integrity is achieved generally in two phases:
@@
    CFI enforcement: This is performed at runtime and involves a reference monitor that checks and enforces the program’s adherence to the CFG derived in the previous step. A common CFI enforcement method is through program instrumentation [1]. Here, a checkpoint $C(p,t)$ is inserted at a control transfer instruction at point $p$ in BBL, with a target, $t$. Here, $p$ and $t$ represent a logical CFG edge between one BBL $v_{i}$ and a target BBL, $v_{j}$. Before executing any control transfer instruction, the program passes control to the reference monitor that verifies whether $C(p,t)$ is associated with $(v_{i},v_{j})\in E$. That is, the monitor ensures that the transfer complies with the CFG. If the check fails, the monitor may halt the program’s execution, throw an exception, or take another enforcement action.
@@ 2.1.1 Control Data Attacks
The primary class are attacks which directly modify control data that influences the processor’s program counter (PC) during execution by modifying return addresses and function pointers [25]. A fundamental attack vector, which we refer to as code injection, involves inserting malicious code into executable memory and directing the processor to execute that code.
@@
Stack-based buffer overflows are a canonical example, which involves overwriting function return addresses and other variables on the stack to redirect execution to code, such as a malicious payload (e.g. shellcode). Heap-based attacks are another approach, involving overflowing buffers on the heap and affecting adjacent heap metadata and objects (e.g. function pointers) that redirect control flow.
@@
A common example of a code-reuse attack is return-oriented programming (ROP) [26], which relies on the manipulation of return instructions to form a sequence of gadgets. Each gadget typically concludes with a return instruction; the next gadget in the sequence starts based on the tampered return address of the previous one. It is known that ROP attacks can achieve Turing-completeness [27]. Jump- (JOP [28]) and call-oriented programming (COP [29, 30]) attacks are two other code reuse attacks, which involve tampering forward control-flow transfers, namely indirect jumps and indirect calls respectively, in order to chain gadgets.
@@ 2.1.2 Non-Control Data Attacks
Rather than directly modifying control data, e.g. function pointers, non-control data attacks rely on the manipulation of more general data to indirectly change a program’s control flow [31, 32, 33]. Detecting non-control-data attacks is challenging because they traverse legitimate edges of the program’s CFG, even though malicious functionality ensues as a result. Let us consider a motivating example. An adversary may execute arbitrary commands on a Linux system if they can influence the arguments to a execve() call in the target program. The target program’s CFG is respected, even though the resulting behaviour is unauthorised. Non-control data attacks can be realised in the following ways:
@@
Branch variable attacks are an example of (A.3), where an adversary influences data used in determining the outcome of conditional branches, i.e. in ‘if’ and ‘switch’ statements. The resulting data causes another legitimate branch to be taken, which nevertheless results in unintended behaviour. [^3] It is important to recognise this semantic gap: just because legitimate edges of a CFG are followed does not mean that unintended or unauthorised behaviour more abstractly cannot be performed. An adversary may also attempt to modify the data used in terminating loops, e.g. ‘while’ and ‘for’ statements, which we refer to as a ‘loop variable attack’.
@@
    Data-oriented attacks: Executes a sequence of existing data gadgets on attacker-controlled inputs to perform malicious actions while respecting a program’s CFG.
@@ 2.2 Control-Flow Transfer Types
In CFI and CFA, monitoring different control-flow transfer types is used to detect particular unauthorised paths taken by a program. It is often undesirable to monitor every possible control-flow transfer due to the significant performance overhead that this can occur, especially on constrained devices. Indeed, not all control-flow transfers are relevant to the security of a system, and it is intuitive to monitor transfers that are common targets for particular exploits of interest. For example, one may wish to pay particular attention to indirect jumps and calls in order to address ROP and JOP attacks. In Figure 2, we give some examples of such transfers used by X86-64 platforms. We now describe the control-flow transfer types which have been used in existing work:
@@
Figure 2: X86-64 assembly examples of control-flow transfers.
@@
    All control-flow transfers: This is the most expansive definition. We assume that $\mathcal{C}$ is a maximal set, or has the greatest cardinality, for the target architecture. In practice, all possible control-flow instructions are examined—conditional jumps, direct calls, indirect calls, etc.—imposing the greatest overhead during monitoring and measurement. It is also the most invasive, requiring the greatest level of instrumentation of the target program (see §6). Monitoring all control-flow transfers has been used in many current proposals [37, 12, 34, 38, 36, 39, 40, 9, 41].
@@
    Conditional direct branches/jumps: Includes all branching and jump instructions that alter the sequence of executed instructions based on a condition. If the condition is true, then the PC is set to a specified address from which it continues execution. In direct jumps, the target address or label is explicitly stated in the instruction. The address is fixed at compile time.
@@
    Return instructions: Used at the end of a function to return to the location of the PC before the function was invoked. On X86, the instruction ret pops the top value from the stack, which should be the return address pushed onto the stack by a previous call instruction, before jumping to that address.
@@ 3 Introducing Control-Flow Attestation
$\mathcal{C}$ denotes the control-flow transfer types (e.g. direct jumps, indirect calls) being monitored, which were presented in §2.2. A general requirement is that a CFA scheme must be able to monitor unique information about a BBL associated with a single control-flow transfer type. During collection, this information may include the addresses of a BBL’s first and final instructions; the control flow transfer’s source and destination addresses, which delineates different BBLs; a unique label identifying the BBL; [^5] IDs are used to overcome issues with Address Space Layout Randomisation (ASLR) in which instructions’ memory addresses change with each program execution. Using only instruction addresses, for example, cannot delineate between different BBLs in this scenario. To solve this challenge, the $T_{\mathcal{P}}$ is instrumented to insert BBL IDs to mark the beginning of BBLs, which do not exist in the original target program. and the type of control-flow transfer and its complete instruction representation, e.g. as machine code.
@@
In Figure 3, we illustrate the knowledge areas that characterise the core elements of CFA schemes. We posit that schemes can be delineated according to their attestation paradigm; the trust anchor; the instrumentation method for collecting measurements; the measurement method; and, finally, the adversarial model under consideration. The coming sections describe the proposals and applications within each area.
@@ 4 Prover-Verifier Paradigms
We recognise that the majority of CFA schemes operate under an interactive paradigm, which is initiated and verified by a (remote) verifier using a challenge-response protocol. This follows a conventional model of remote attestation used by trusted computing systems, such as the Trusted Platform Module (TPM) [48]. However, other approaches have also been developed in recent years.
4.1 Remote Attestation
@@ 4.2 Continuous Reporting
Rather than reporting measurements in a challenge-response protocol, LiteHAX by Dessouky et al. [34] reports measurements continuously $\mathcal{P}$to $\mathcal{V}$. In this work, there is an initial offline phase where $\mathcal{V}$ generates the CFG of $T_{\mathcal{P}}$. However, during the online phase, $\mathcal{P}$ sends attestation reports to $\mathcal{V}$ in a one-way, continuous manner. There is no formal challenge-response protocol at run-time. We illustrate this in Figure 5. ACFA [44] similarly reports measurements $\mathcal{P}$to $\mathcal{V}$. The transmission of the attestation report is triggered by one of three conditions: (1) the expiration of a timer, which enables $\mathcal{P}$ to send reports periodically; (2) the memory designated for storing measurement results becomes full; or (3) the completion of the attested code execution. Another difference with ACFA is that, after sending the attestation report, $\mathcal{P}$ pauses the target program’s execution until $\mathcal{V}$ returns a response to notify whether execution can be resumed.
Figure 5: LightHAX: a uni-directional CFA scheme [34].
@@ 4.3 Hybrid Attestation
We define hybrid attestation as a combination of local monitoring and remote attestation. Geden and Rasmussen [47] follow the same offline stage as conventional remote attestation where $\mathcal{V}$ extracts a CFG, known as a runtime integrity model, before deploying the proving device. The model is stored internally in a discrete hardware security module (HSM) aboard $\mathcal{P}$. The HSM continuously takes control-flow events as input from the CPU in real-time in order to thwart time-of-check-to-time-of-use (TOCTOU) problem in conventional remote attestation. A remote verifier can acquire a report from the HSM, showing whether any security policy violations were detected over the course of the target program’s execution.
@@
GuaranTEE by Morbitzer et al. [10] investigates control-flow attestation in untrusted cloud environments. The work differs in that $\mathcal{P}$ and a local verifier, $\mathcal{V}^{\prime}$, run on the same cloud-based virtual machine (VM) in Intel SGX enclaves. $\mathcal{P}$ is assumed to host a security-sensitive application that may be targeted by control-flow attackers; a signing service for Covid-19 digital passports is given as an example. $\mathcal{V}^{\prime}$ measures and analyses valid control-flow transfers, which are passed through a trampoline implemented in $\mathcal{P}$. We illustrate this in Figure 6. The target program’s instrumentation is performed through extensions to the LLVM compiler framework, allowing $\mathcal{P}$’s full CFG to be constructed pre-deployment. GuaranTEE supports a form of remote and local attestation. At launch-time, each SGX enclave mutually attests the other on the local machine, while enabling a remote $\mathcal{V}$ to retrieve attestation reports from $\mathcal{V}^{\prime}$.
@@ 4.5 Collective Attestation
Figure 7: The ARCADIS system model [41].
@@
Li et al. [50] proposed CFRV for attesting networks of peer-to-peer IoT devices, where devices may act as $\mathcal{P}$ and $\mathcal{V}$. The authors address the problem of adversarial verifiers who submit malicious attestation challenges that purposely induce a particular control-flow path resulting in unintended behaviour. To address this, the proposal combines secret sharing and mutual attestation to create an agreed attestation measurement between multiple devices.
@@
Figure 8: System overview of DIAT [45]. Unlike conventional schemes, measurements from each sub-system’s CFGs are aggregated within attestation reports.